Systems Staff Meeting Notes
June 25, 2007
June 25, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Review Minutes of 6/18/07
- "Ocean's 13" was funny.
- Monday, July 31, the new Dean of the Libraries will be attending the Systems Department meeting.
- List of Systems departmental accomplishments to be shared at the West convocation - send Will anything you feel should be mentioned.
Ongoing Systems Updates
- Room 7 - DLC has a need for it. We are currently storing a lot of the old DLC monitors in there. Discussion ensues on the distribution of machines for public use. Amy talked to Laurie Brennan about securing CPU's in MSL. We will order cradles or cables for security, based on cost. We are letting the Directors make the decision.
- Nets #6 - delete.
- Orders on hold until July 10.
- SAN training for Logan - Bill has approved it. If we have to pay in advance to register, we'll have to try and reimburse.
- VM Ware training has also been approved.
- Vista training on hold. Image-pushing upgrade? Logan reports that it breaks IRDB & PAULA. May have to do it as a fresh install. He will set it up correctly, and then create the image.
- Campus VISTA images came down with Office 2007 - Logan is testing it.
- Staff training - issue with ALEPH v.18 conflict? It works now with VISTA, but it has not been fully tested.
- Public computer replacements - we're working on it. Security for the 270's is an issue. Once we have a method to secure them and know what needs to go where, we can get them out. Amy has a meeting scheduled with Evan regarding the preferred method of upgrading RAM.
- Imaging DHCP - solution is August.
Software Projects Status
- Xerox - 1st PDF to print problem still exists. Amy has received two SYSHELPS about the print queue being backed up - Xerox was contacted to resolve the issue.
- Exchange mobile device synch problem is still there - Will is working on it.
- Campus AD - no news.
- Data transfer routines for Grover - should be done tomorrow or Wednesday. We need to schedule another group meeting to discuss the ways this new application will affect the way business is done.
- Tracking database - nothing has been released to Erich yet, but it will be with accepted testing.
- Mark has been working on Ephemeral Cities. He had a presentation on Sunday and worked all weekend. Laurie mentioned at the Division meeting today that the DLC was pleased with the effort Systems made to get the servers in better working order. The RAM upgrade seems to have done the trick.
- Blog testing ongoing.
- Patron Load file sent to Jennifer Koontz at FCLA with the changes for ALEPH 18. They want us to encrypt all traffic between us and FCLA. They also want to take over sending Accounts Payable files to UF.
- Lori Driscoll wants Patron Load File changed for the summer to enable students accepted in the fall to have borrowing privileges.
Systems Security Management
- Metalib v. 4 - Tom is only on the project as a consultant. Michael Jay, Jason Fleming and Rich are handling it.
- Link in ALEPH catalog now goes to Endeca.
- Link on homepage to new Endeca.
- New homepage design to go into effect on Wednesday.
- Tom is off all next week. He needs to meet with his backup this week for instruction.
- Meeting held.